OpenStudy (anonymous):

A triangle has a leg that measures 3 and a hypotenuse of 9. Which leg measure would form a right triangle?

OpenStudy (anonymous):

\[ 6 \sqrt{2} \]

OpenStudy (anonymous):

these are the options

OpenStudy (anonymous):

what is \( 36 \times 2 \) ? That's correct it is 72. So \( \sqrt{72} \) is the right answer.

OpenStudy (anonymous):

How did you find it out? Where did the 36 come from

OpenStudy (anonymous):